### v1

@@ 4 Conclusions
Conducting eigenvalue-based timescale analysis on the KS equations proved instrumental in isolating the impact of individual training hyperparameters on timescale elimination in the latent space. Notably, the number of convolutional layers in the autoencoder and the latent dimensionality exhibited little discernible effect on latent space timescales. Interestingly, the training methodology — in terms of coupled versus de-coupled optimization of the autoencoder and neural ODE components — also did not have any effect on the time-scale reduction, although, as reported in recent work Wan2023 , the decoupled training approach had lower single-step and rollout errors for both PDE systems tested here.
@@
Efforts to establish a framework for a priori identification of this optimal $n_{t}$ value based on the problem’s physics involved analyzing the largest timescales ($t_{max}$) in the latent space for various $n_{t}$ values. For the KS results observed here, the largest time-scales of the neural ODE exhibiting the best predictive accuracy were found to align closely with those of the full-order system, pointing to the role of capturing slow-moving dynamics in the surrogate latent space for enhanced accuracy. For the unsteady detonation, a similar general trend was seen in that the rollout error was shown to be minimized at an particular value of $n_{t}$. Although promising, further studies need to be done to investigate both the generality of time-scale relationships observed here and neural ODE feasibility in modeling the compressible reacting Navier-Stokes equations. Such aspects will be explored in future work.
Acknowledgements
This research used resources of the Argonne Leadership Computing Facility (ALCF), a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357. The authors would also like to acknowledge the ALCF Summer intern program. SB acknowledges support from the AETS fellowship from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy under contract DE-AC02-06CH11357. RM acknowledges funding support from ASCR for DOE-FOA-2493 “Data-intensive scientific machine learning”.

### v3

@@ 4 Conclusions
Conducting eigenvalue-based timescale analysis on the KS equations proved instrumental in isolating the impact of individual training hyperparameters on timescale elimination in the latent space. Notably, the number of convolutional layers in the autoencoder and the latent dimensionality exhibited little discernible effect on latent space timescales. Interestingly, the training methodology — in terms of coupled versus de-coupled optimization of the autoencoder and neural ODE components — also did not have any effect on the time-scale reduction, although, as reported in recent work [55], the decoupled training approach had lower single-step and rollout errors for both PDE systems tested here.
@@
Efforts to establish a framework for a priori identification of this optimal $n_{t}$ value based on the problem’s physics involved analyzing the largest timescales ($t_{max}$) in the latent space for various $n_{t}$ values. For the KS results observed here, the largest time-scales of the neural ODE exhibiting the best predictive accuracy were found to align closely with those of the full-order system, pointing to the role of capturing slow-moving dynamics in the surrogate latent space for enhanced accuracy. For the unsteady detonation and atmospheric flow datasets, a similar general trend was seen in that the rollout error was shown to be minimized at an particular value of $n_{t}$. Although promising, further studies need to be done to investigate both the generality of time-scale relationships observed here and neural ODE feasibility in modeling the compressible reacting Navier-Stokes equations and atmospheric flows. Additionally, benchmarking different surrogate modeling architectures (such as DeepONets [89], Fourier neural operators [90], and other approaches), and assessing the transferability of the latent time scale trends uncovered in this work to these different architectures, is a useful direction. Such aspects will be explored in future work.
Acknowledgments
This research used resources of the Argonne Leadership Computing Facility (ALCF), a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357. The authors would also like to acknowledge the ALCF Summer intern program. SB and PP acknowledge laboratory-directed research and development (LDRD) funding support from Argonne’s Advanced Energy Technologies (AET) directorate through the Advanced Energy Technology and Security (AETS) Fellowship. RM acknowledges funding support from ASCR for DOE-FOA-2493 “Data-intensive scientific machine learning”.